OpenStudy (anonymous):

Using differentiation, find the equation of the tangent to the curve y=4+ln(x+1) at the point where x=0

sam (.sam.):

dy/dx= 1/(x+1)

sam (.sam.):

when x=0, dy/dx= 1/1 dy/dx=1

OpenStudy (anonymous):

Yes. I got that. But I don't know how to get the equation?

sam (.sam.):

you need to find the coordinates

sam (.sam.):

using x=0, substitute into original y=4+ln(x+1) then y=4

sam (.sam.):

you have (0, 4) and gradient, m=1

sam (.sam.):

\[y-y _{1}=m(x-x _{1})\]

OpenStudy (anonymous):

Oh, Ok. so the answer is y=x+4 :)

sam (.sam.):

yep

OpenStudy (anonymous):

Thank you! :)

sam (.sam.):

no problem :)